If a dog has a mass of 20.5 kg, what is its mass in the following units? Use scientific notation in all of your answers. (a) What is the dog's mass in grams? × 10 g (b) What is the dog's mass in milligrams? × 10 mg (c) What is the dog's mass in micrograms? × 10 μ g
Answer #1

a)

Conversion to be used:

1 Kg = 1000 g

mass = 20.5 Kg

= 20.5 * 1000 g

= 2.05*10^4 g

Answer: 2.05*10^4 g

b)

Conversion to be used:

1 Kg = 1000 g

1 g = 1000 mg

mass = 20.5 Kg

= 2.05*10^4 g

= 2.05*10^4 * 1000 mg

= 2.05*10^7 mg

Answer: 2.05*10^7 mg

c)

Conversion to be used:

1 Kg = 1000 g

1 g = 1.0*10^6 ug

mass = 20.5 Kg

= 2.05*10^4 g

= 2.05*10^4 * 1.0*10^6 ug

= 2.05*10^10 ug

Answer: 2.05*10^10 ug